Monthly spending ($)

 

 

 

 

 

 

Age

20s

30s

$0.00 - $49.99

75

29

$50.00 - $99.99

91

83

$100.00 - $149.99

113

90

$150.00 - $199.99

89

128

$200.00 - $249.99

63

159

$250.00 or more

34

46

Total customers

465

535

 

a. What is the probability that a customer spends at least $150.00 or is in his/her 30s?



1
Expert's answer
2021-05-18T08:03:03-0400

The table of this question will be:

Ages

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Monthly\ spending(\$) & 20s & 30s \\ \hline
 0.00\$-49.99\$ & 75 & 29\\
50.00\$-99.99\$ & 91& 83\\
100.00\$-149.99\$ & 113 & 90 \\
150.00\$-199.99\$ & 89 & 128\\
200.00\$-249.99\$ & 63 & 159\\
250.00\$ \ or\ more & 34 & 46\\
 \hdashline
 \text{Total customers} & 465 & 535
\end{array}"

The probability that a customer spends at least $150.00 or is in his/her 30s will be


"P\big((\geq150.00\$)\cup(30s)\big)"

"=\dfrac{535}{465+535}+\dfrac{89+63+34}{465+535}=0.721"

The probability that a customer spends at least $150.00 or is in his/her 30s is "0.721."